solve the following logarithmic equation
log(x-5)=2
5+log2x=6

Answer by stanbon(48535)

solve the following logarithmic equation
log(x-5)=2
x-5 = 10^2
x-5 = 100
x = 105
-------------------
5+log2x=6
log(2x) = 1
2x = 10^1
x = 5
=====================
Cheers,
Stan H.